# STONE: A Submodular Optimization Framework for Active 3D Object Detection

### v1

@@ 5.2 Results
KITTI Dataset Results. We evaluate the performance of STONE against other baseline methods in Table 1. We clearly notice that STONE outperforms all the prior active learning methods, irrespective of the detection difficulty level. In particular, on average, we observe 3D AP improvements of 1.47%, 0.84%, and 1.24%, over CRB [34] for the EASY, MODERATE, and HARD evaluation modes. We also observe improvements of 0.54%, 0.08%, and 0.63% over KECOR [33] respectively. The BEV scores are available in the appendix for reference.
From Figure 2, it is evident that regardless of the detection difficulty level, STONE consistently surpasses other baseline methods. Specifically, compared to the previous state-of-the-art methods KECOR [33] and CRB [34], with 1% of labeled bounding boxes from the entire pool of unlabeled point clouds at the HARD difficulty level, STONE is on average 0.9% higher than KECOR [33] and 1.9% higher than CRB [34]. Although STONE requires 37% more annotated bounding boxes than KECOR [33], it achieves higher accuracy in earlier query rounds, resulting in faster training times due to fewer epochs and active queries needed. Compared to CRB [34], even with 36% more labeled bounding boxes at the final round of active queries, STONE is still on average 0.84% higher across all the difficulty levels.
Table 1: 3D AP(%) scores on KITTI [15] validation set with 1% queried bounding boxes, using PV-RCNN [46] as the backbone detection model.
@@
Figure 2: 3D mAP (%) of AL baselines on the KITTI [15] validation set with PV-RCNN [46].
Figure 3: 3D mAP (%) of AL baselines on the Waymo Open [49] validation set with PV-RCNN [46].
Waymo Open Dataset Results. To further test the generality and robustness of STONE, we also evaluate our approach on the Waymo Open dataset [49] using the APH score as the performance metric across the two levels of difficulty LEVEL 1 and LEVEL 2. As shown in Figure 3, STONE surpasses other baseline methods as the number of bounding boxes increases. Compared to previous state-of-the-art methods KECOR [33] and CRB [34] at 25,000 bounding boxes, STONE achieves higher accuracy for both levels of detection difficulty. Our method uses fewer active learning loops than KECOR [33], meaning it reaches a 2% higher APH score with fewer training epochs. Additionally, when comparing STONE with CRB [34] in the final active learning round, STONE achieves a 1% higher APH score on average with 24% fewer bounding boxes.
@@
2D Object Detection Results. We also demonstrate the generalizability of our method on a 2D object detection task. We perform experiments on the PASCAL VOC [9] dataset, that contains 20 object classes. For the experiments, we use VOC07 trainval and VOC07+12 trainval to train a Single Shot MultiBox Detector (SSD) [32] with a VGG-16 backbone [47], and test on VOC07 test. We follow the training guidelines and setup as outlined in AL-MDN [5] to reproduce the results. We set $N_{q}$ to be 1000 for all the methods, for a total of three queries. In Table 2, we summarize the results in terms of mean Average Precision (mAP), noting that our method either outperforms or is comparable to all baseline methods. In the first query, we observe improvements of 2.93% and 2.43% over AL-MDNgmm and AL-MDNeff respectively. Our method, while achieving a balance in label distribution, also scales well to large datasets with more semantic labels.
Computational Complexity. STONE achieves significant GPU memory savings compared to KECOR. On the KITTI dataset [15], with a batch size of 6 (for a fair comparison), our method consumes 10 GB of GPU memory, whereas KECOR [33] consumes 24 GB, which is 140% more GPU memory. Additionally, STONE maintains similar running time to KECOR [33].
6 Ablation Study
We perform ablation studies, on KITTI [15], to assess the efficacy of our approach and to better understand the key mechanisms and components involved. In the ablation study, all the experiments use 5 rounds of active learning selection, acquiring a total of 500 point clouds for annotation. We present and discuss additional ablation experiments in the Appendix (8).
Contributions of $\hat{L}_{reg}$ and $\hat{L}_{cls}$. Given that our method incorporates both regression loss $\hat{L}_{reg}$ and classification loss $\hat{L}_{cls}$ to handle potential class imbalance, we have carried out extensive experiments to analyze their impact on model performance. We first investigated whether regression or classification loss has a greater impact by using gradients generated from each loss separately during the active learning stage. The results showed that using only $\hat{L}_{reg}$ resulted in an average 3D AP drop of 0.77% for HARD while using only $\hat{L}_{cls}$ resulted in an average 3D AP drop of 3.12% for HARD on the KITTI [15] validation dataset. This indicates that $\hat{L}_{cls}$ has a larger impact on our model’s performance, and both losses are essential for optimal results.
Effect of the reweighing factor. To further evaluate the reweighing factor’s importance on our method’s performance, we perform three crucial experiments. In Table 6(row 1), we study the effect of not having the reweighing factor on both the regression and classification losses. We observe performance drops of 3.78%, 1.55%, and 2.45% for the EASY, MODERATE, and HARD levels of difficulty, respectively. We also observe similar performance drops in Table 6(row 2) and Table 6(row 3) with no reweighing factor for the regression or classification loss, respectively. This fundamental issue arises because, in contrast to classification tasks where a single loss function is typically considered, detection tasks require the simultaneous optimization of multiple loss components. Specifically, reweighing only the classification or regression loss in detection tasks can disrupt the balance within the model, as it does not adequately model the interdependencies between these loss functions, leading to the suboptimal performance of the detector.
@@ 6 Ablation Study
Relevance of each stage. In Table 6, we examine the importance and the relevance of each stage of our method. We assess the performance using 3D AP (%) and the number of bounding boxes annotated with 500 point cloud selections. From Table 6 we observe that the removal of any single component leads to a drop in 3D AP performance. In particular, using the GBMSS stage only leads to a drop of 3D AP by 3.43%, 3.83%, and 2.55% for the EASY, MODERATE and HARD difficulty levels respectively. In such a scenario, the bounding boxes annotated increase by 71.43%, which shoots up the labeling cost. It is interesting to note that while using only the SDMCB stage results in annotating the lowest bounding boxes, it also leads to a drop in performance. In conclusion, both stages of our method are necessary for the optimal selection of samples, with a lower labeling cost, to ensure the efficiency and accuracy of the proposed active learning method.
@@ 7 Conclusion
In this paper, we propose a novel approach called STONE, a unified active 3D object detection methodology, based on submodular optimization. We provide a robust and compute-efficient solution by proposing a two-stage algorithm that first utilizes a submodular function based on the gradients using the Monte Carlo dropout [13] in order to select representative point clouds. We then apply a greedy search algorithm to balance the data distribution due to the possibility of having an imbalance across all the labeled point clouds. Extensive experiments on benchmark autonomous driving datasets, including KITTI [15] and Waymo Open [49], demonstrate the effectiveness and generalizability of our method.
Limitations. One limitation of the proposed method is that it does not reduce the running time of existing active 3D detection methods, which is primarily due to the large number of unlabeled point clouds. As future work, we will investigate more efficient methods for active 3D detection.
@@ 54. Answer: [N/A]
8 Appendix / supplemental material
8.1 More ablation studies
Table 5: Impact of different submodular functions.
3D AP in % Bounding Boxes
Submodular functions EASY MOD. HARD
Facility location 82.91 70.64 65.98 3121
Max coverage 81.46 69.47 64.68 1674
STONE 83.03 70.33 65.33 1530
Table 6: Sensitivity to thresholds $\Gamma_{1}$, $\Gamma_{2}$.
3D AP in %
Components EASY MOD. HARD
$\Gamma_{1}$: 400, $\Gamma_{2}$: 200 82.32 68.24 64.08
$\Gamma_{1}$: 400, $\Gamma_{2}$: 300 83.03 70.33 65.33
$\Gamma_{1}$: 500, $\Gamma_{2}$: 200 78.62 67.44 63.77
$\Gamma_{1}$: 500, $\Gamma_{2}$: 200 80.70 69.18 64.01
$\Gamma_{1}$: 600, $\Gamma_{2}$: 200 82.53 68.90 63.82
$\Gamma_{1}$: 600, $\Gamma_{2}$: 300 81.67 67.47 64.72
Impact of different submodular functions. We evaluate STONE with different submodular functions used in stage 1 i.e., GBMSS, and report the results in Table 8.1. Submodular functions such as facility location have demonstrated effectiveness in modeling representativity and diversity [22], as well as in active learning [27, 26]. We employ these functions in our GBMSS study to assess their effectiveness. In terms of performance measured by 3D AP, we observe results similar to those of STONE. However, it comes with an extremely high labeling cost with a 50.98% rise in bounding box annotations. With the max coverage submodular function, we see slight performance drops of 1.57%, 0.06%, and 0.65% for the EASY, MODERATE and HARD difficulty levels respectively. Overall, our method attains the best results in terms of 3D average precision (AP) while maintaining a low labeling budget.
Sensitivity to thresholds $\Gamma_{1}$ and $\Gamma_{2}$. We conduct a sensitivity analysis of model performance to different variants of thresholds $\Gamma_{1}$ and $\Gamma_{2}$, at all difficulty levels, to understand their importance in our method. We report the 3D AP (in %) across six different combinations in Table 8.1 with 500 point-cloud selections. For the MODERATE and HARD difficulty levels, we observe fluctuations within 2.89% and 1.56% (compared to the lowest). This suggests that our method is relatively stable with the best $\Gamma_{1}$ and $\Gamma_{2}$ being 400 and 300 respectively.
Figure 4: STONE: Pipeline of Submodular Optimization Framework for Active 3D Object Detection.
Figure 5: Illustration of the entropy after step 1 and step 2 of SDMCB to 10 active learning query rounds.

### v2

@@ 5.2 Results
KITTI Dataset Results. We evaluate the performance of STONE against other baseline methods in Table 1. We clearly notice that STONE outperforms all the prior active learning methods, irrespective of the detection difficulty level and backbone model. In particular with PV-RCNN [49] as backbone, on average, we observe 3D AP improvements of 1.47%, 0.84%, and 1.24%, over CRB for the EASY, MODERATE, and HARD evaluation modes. We observe improvements of 0.54%, 0.08%, and 0.63% over KECOR respectively.
From Figure 3, it is evident that regardless of the detection difficulty level, STONE consistently surpasses other baseline methods. Specifically, compared to the previous state-of-the-art methods KECOR and CRB, with 1% of labeled bounding boxes from the entire pool of unlabeled point clouds at the HARD difficulty level, STONE is on average 0.9% higher than KECOR and 1.9% higher than CRB. Although STONE requires 37% more annotated bounding boxes than KECOR, it achieves higher accuracy in earlier query rounds, resulting in faster training times due to fewer epochs and active queries needed. Compared to CRB, even with 36% more labeled bounding boxes at the final round of active queries, STONE is still on average 0.84% higher across all the difficulty levels.
Table 1: 3D AP(%) scores on KITTI validation set with 1% queried bounding boxes, using PV-RCNN as the backbone detection model.
@@
Figure 3: 3D mAP (%) of AL baselines on the KITTI validation set with PV-RCNN.
Figure 4: 3D mAP (%) of AL baselines on the Waymo Open validation set with PV-RCNN.
Waymo Open Dataset Results. To further test the generality and robustness of STONE, we also evaluate our approach on the Waymo Open dataset using the APH score as the performance metric across the two levels of difficulty LEVEL 1 and LEVEL 2. As shown in Figure 4, STONE surpasses other baseline methods as the number of bounding boxes increases. Compared to previous state-of-the-art methods KECOR and CRB at 25,000 bounding boxes, STONE achieves higher accuracy for both levels of detection difficulty. Our method uses fewer active learning loops than KECOR, meaning it reaches a 2% higher APH score with fewer training epochs. Additionally, when comparing STONE with CRB in the final active learning round, STONE achieves a 1% higher APH score on average with 24% fewer bounding boxes.
2D Object Detection Results. We also demonstrate the generalizability of our method on a 2D object detection task. We perform experiments on the PASCAL VOC [9] dataset, that contains 20 object classes. For the experiments, we use VOC07 trainval and VOC07+12 trainval to train a Single Shot MultiBox Detector (SSD) [33] with a VGG-16 backbone [50], and test on VOC07 test. We follow the training guidelines and setup as outlined in AL-MDN [5] to reproduce the results. For all methods, $N_{q}$ is set to be 1000 for a total of three queries. In Table 2, we summarize the results in terms of mean Average Precision (mAP), noting that our method either outperforms or is comparable to all baseline methods. In the first query, we observe improvements of 2.93% and 2.43% over AL-MDNgmm and AL-MDNeff respectively. Our method, while achieving a balance in label distribution, also scales well to large datasets with more semantic labels.
@@
Computational Complexity. STONE achieves significant GPU memory savings compared to KECOR, which computes gradients of the output of the ROI head’s fully connected shared layer, resulting in a gradient matrix of high dimensions and memory. In contrast, STONE focuses gradient computation on the outputs of the classification and regression loss layers within the ROI head, which are much lower in dimensionality. On the KITTI dataset, with a batch size of 6 (for a fair comparison), our method consumes 10 GB of GPU memory, whereas KECOR consumes 24 GB, which is 140% more GPU memory. Additionally, STONE maintains a similar running time to KECOR.
@@
6 Ablation Study
We perform ablation studies, on KITTI, to assess the efficacy of our approach and to better understand the key mechanisms and components involved. In the ablation study, all the experiments use 5 rounds of active learning selection, acquiring a total of 500 point clouds for annotation. We present and discuss additional ablation experiments in the Appendix (8).
Contributions of $\hat{L}_{reg}$ and $\hat{L}_{cls}$. Given that our method incorporates both regression loss $\hat{L}_{reg}$ and classification loss $\hat{L}_{cls}$ to handle potential class imbalance, we have carried out extensive experiments to analyze their impact on model performance. We first investigated whether regression or classification loss has a greater impact by using gradients generated from each loss separately during the active learning stage. The results showed that using only $\hat{L}_{reg}$ resulted in an average 3D AP drop of 0.77% for HARD while using only $\hat{L}_{cls}$ resulted in an average 3D AP drop of 3.12% for HARD on the KITTI validation dataset. This indicates that $\hat{L}_{cls}$ has a larger impact on our model’s performance, and both losses are essential for optimal results.
Effect of the reweighing factor. To further evaluate the reweighing factor’s importance on our method’s performance, we perform three crucial experiments. In Table 5.2 (row 1), we study the effect of not having the reweighing factor on both the regression and classification losses. We observe performance drops of 3.78%, 1.55%, and 2.45% for the EASY, MODERATE, and HARD levels of difficulty, respectively. We also observe similar performance drops in Table 5.2 (row 2) and Table 5.2 (row 3) with no reweighing factor for the regression or classification loss, respectively. This fundamental issue arises because, in contrast to classification tasks where a single loss function is typically considered, detection tasks require the simultaneous optimization of multiple loss components. Specifically, reweighing only the classification or regression loss in detection tasks can disrupt the balance within the model, as it does not adequately model the interdependencies between these loss functions, leading to the suboptimal performance of the detector.
Relevance of each stage. In Table 5.2, we examine the importance and the relevance of each stage of our method. We assess the performance using 3D AP (%) and the number of bounding boxes annotated with 500 point cloud selections. From Table 5.2 we observe that the removal of any single component leads to a drop in 3D AP performance. In particular, using the GBMSS stage only leads to a drop of 3D AP by 3.43%, 3.83%, and 2.55% for the EASY, MODERATE and HARD difficulty levels respectively. In such a scenario, the bounding boxes annotated increase by 71.43%, which shoots up the labeling cost. It is interesting to note that while using only the SDMCB stage results in annotating the lowest bounding boxes, it also leads to a drop in performance. In conclusion, both stages of our method are necessary for the optimal selection of samples, with a lower labeling cost, to ensure the efficiency and accuracy of the proposed active learning method.
Backbone-agnostic performance. In all of the experiments in this paper, we use PV-RCNN as the backbone detection model. However, to show the invariance of our method towards a change in the backbone, we perform experiments with SECOND [61], a widely used 3D object detector. The results, as shown in Table 5, indicate that STONE achieves a 3.4% higher 3D mAP score at the HARD level and 2.43% higher mAP score at the HARD level in BEV detection compared to the state-of-the-art method, KECOR. This demonstrates the performance and generality of the proposed approach.
Table 5: 3D mAP(%) scores on KITTI validation set with 1% queried bounding boxes, using SECOND [61] as the backbone detection model.
3D Detection average mAP BEV Detection average mAP
Method EASY MOD. HARD EASY MOD. HARD
Random 66.33 55.48 51.53 75.66 63.77 59.71
CORESET [44] 66.86 53.22 48.97 73.08 61.03 56.95
LLAL [62] 69.19 55.38 50.85 76.52 63.25 59.07
BADGE [1] 69.92 55.60 51.23 76.07 63.39 59.47
BAIT [21] 69.45 55.61 51.25 76.04 63.49 53.40
CRB [35] 72.33 58.06 53.09 78.84 65.82 61.25
KECOR [34] 74.05 60.38 55.34 80.00 68.20 63.20
STONE 76.86 64.04 58.75 82.14 70.82 65.68
@@ 7 Conclusion
In this paper, we propose a novel approach called STONE, a unified active 3D object detection methodology, based on submodular optimization. We provide a robust and compute-efficient solution by proposing a two-stage algorithm that first utilizes a submodular function based on the gradients using the Monte Carlo dropout [13] to select representative point clouds. We then apply a greedy search algorithm to balance the data distribution due to the possibility of having an imbalance across all the labeled point clouds. Extensive experiments on benchmark autonomous driving datasets, including KITTI [15] and Waymo Open [52] datasets, demonstrate the effectiveness and generalization of our method.
Limitations. One limitation of the proposed method is that it does not reduce the running time of existing active 3D detection methods, which is primarily due to the large number of unlabeled point clouds. In future work, we will investigate more efficient methods for active 3D object detection.
@@
Acknowledgements. We would like to thank the anonymous reviewers for their helpful comments. This project was supported by a grant from the University of Texas at Dallas. This work is also supported by the National Science Foundation under Grant Numbers IIS-2106937, a gift from Google Research, an Amazon Research Award, and the Adobe Data Science Research award.